The points A,B,C and D are consecutive vertices of a regular polygons with 20 sides.Calculate angles
(a) ABC
(b) ACD
(c) ABD

Answer by greenestamps(13200)



ABC is an interior angle of the regular 20-sided polygon. The measure of an exterior angle is 360/20 = 18 degrees, so the measure of an interior angle is 180-18 = 162 degrees.

ANSWER (a): ABC = 162 degrees

Triangle ABC is isosceles, and angle ABC is 162 degrees. That makes angles BAC and BCA each 9 degrees. Angle ACD is angle BCD minus angle BCA, which is 162-9 = 153 degrees. And angle ABD is the same size angle.

ANSWERS (b) and (c): ABD = ACD = 153 degrees